OpenStudy (sphott51):

solve 125^9x-2 = 150 in log ir exponentials

OpenStudy (anonymous):

firstly, do you know the log properities?

OpenStudy (sphott51):

Uhh like log 125^9x-2 = log150

OpenStudy (anonymous):

if you take the log of a^b what happens?

OpenStudy (sphott51):

(9x-2)log125=log150

OpenStudy (anonymous):

the b comes down and multiplies with the rest. i.e. you will get b*log(a)

OpenStudy (anonymous):

ok what do you think you should be next?

OpenStudy (sphott51):

Uhh... (9x)log125 - (2)log125 = log 150 ?

OpenStudy (anonymous):

no divide out log125

OpenStudy (anonymous):

then simply just solve for x.

OpenStudy (sphott51):

Uhh how do I do that.